OpenStudy (anonymous):

can you explain???? What is the solution to -2(3x - 9) + 5x = -10? -28 -8 8 28

OpenStudy (nikato):

Distribute to get -6x+18+5x=-10

OpenStudy (anonymous):

okay I got that first part right.

OpenStudy (nikato):

Add like terms -x+18=-10

OpenStudy (nikato):

Subtract 18 from both sides And finally multiply both sides by -1

OpenStudy (anonymous):

okay im lost??? @nikato

OpenStudy (nikato):

Ok subtract 18 from both sides -x+18=-10 -18. -18 --------------- -x=-28 Multiply both sides by -1 x=28

OpenStudy (anonymous):

how did u get rid of -6x and 5x

OpenStudy (nikato):

Combine like terms -6x+5x=-x It's the same thing without the x. -6+5=-1